OpenStudy (anonymous):

Either use factoring or the quadratic formula to solve the given equation. (log10 x)^2 + log10 x = 6

OpenStudy (jdoe0001):

\(\bf (log_{10}x)^2+log_{10}x=6 ?\)

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle {\rm let~}~~~~a= (\log_{10}~x) }\) just to visualize it.

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle a^2+a=6 }\)

OpenStudy (solomonzelman):

then subtract 6 from both sides, and factor

OpenStudy (anonymous):

so you would get 3 and -2

OpenStudy (solomonzelman):

a^2 + a = 6 a^2 + a - 6 = 0 (a + 3)(a - 2) = 0 a=-3 or 2

OpenStudy (solomonzelman):

(( you were almost right ))

OpenStudy (solomonzelman):

So, \(\large\color{black}{ \displaystyle \log_{10}x=-3 }\) and \(\large\color{black}{ \displaystyle \log_{10}x=2 }\)

OpenStudy (solomonzelman):

which one does not work ?

OpenStudy (anonymous):

-3

OpenStudy (solomonzelman):

yup

OpenStudy (solomonzelman):

So, all you are left with is: \(\large\color{black}{ \displaystyle \log_{10}x=2 }\)

OpenStudy (anonymous):

so the answers would be log10x=2...can you simplify that at all?

OpenStudy (solomonzelman):

yes, you can solve for x.

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \log_{A}B=C~~~~~\Longrightarrow~~~~~~A^C=B }\)

OpenStudy (anonymous):

so 10^3

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \log_{10}x=2~~~~~\Longrightarrow~~~~~~~~x=? }\)

OpenStudy (solomonzelman):

retry

OpenStudy (solomonzelman):

(or typo)

OpenStudy (anonymous):

10^2

OpenStudy (solomonzelman):

yup

OpenStudy (anonymous):

so 1000 is my answer

OpenStudy (solomonzelman):

10^2=1000 ?

OpenStudy (solomonzelman):

are you sure ?

OpenStudy (anonymous):

100 sorry typo

OpenStudy (solomonzelman):

yes, I can show how to make the keyboard slower (less "sensitive")....

OpenStudy (solomonzelman):

anyway.... x=100 (that is your only solution that works )

OpenStudy (anonymous):

my homework says it doesnt work...hmmmm

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle (\log_{10}x)^2+(\log_{10}x)=6 }\) \(\large\color{black}{ \displaystyle (\log_{10}100)^2+(\log_{10}100)=6 }\) \(\large\color{black}{ \displaystyle (\log_{10}10^2)^2+(\log_{10}10^2)=6 }\) \(\large\color{black}{ \displaystyle (2\log_{10}10)^2+(2\log_{10}10)=6 }\) \(\large\color{black}{ \displaystyle (2)^2+(2)=6 }\) \(\large\color{black}{ \displaystyle 4+2=6 }\)

OpenStudy (solomonzelman):

I don't see any fault in x being 100. and it is 100....

OpenStudy (solomonzelman):

tell your teachers that THEY don't work !!!

OpenStudy (solomonzelman):

wait, I think there is another answer...

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle (\log_{10}x)=-3 }\) \(\large\color{black}{ \displaystyle x=10^{-3}=1/1000 }\)

OpenStudy (solomonzelman):

yes, that one can also work.... ((you can not take a log of a negative number, but it can be equal to a negative number))

OpenStudy (solomonzelman):

So, 2 solutions x=1/100 x=1/1000